### v1

# TopoCellGen: Generating Histopathology Cell Topology with a Diffusion Model

Abstract
Accurately modeling multi-class cell topology is crucial in digital pathology, as it provides critical insights into tissue structure and pathology. The synthetic generation of cell topology enables realistic simulations of complex tissue environments, enhances downstream tasks by augmenting training data, aligns more closely with pathologists’ domain knowledge, and offers new opportunities for controlling and generalizing the tumor microenvironment. In this paper, we propose a novel approach that integrates topological constraints into a diffusion model to improve the generation of realistic, contextually accurate cell topologies. Our method refines the simulation of cell distributions and interactions, increasing the precision and interpretability of results in downstream tasks such as cell detection and classification. To assess the topological fidelity of generated layouts, we introduce a new metric, Topological Fréchet Distance (TopoFD), which overcomes the limitations of traditional metrics like FID in evaluating topological structure. Experimental results demonstrate the effectiveness of our approach in generating multi-class cell layouts that capture intricate topological relationships.
1 Introduction
Deep-learning methods have made substantial advances in foundational tasks for nuclei analysis, including instance segmentation [29, 21, 40, 32], classification, and detection [1, 52]. These tasks provide the basis for downstream analyses, enabling detailed characterization of tissue architecture and cellular interactions, which are crucial for diagnostic and prognostic applications in pathology [46, 12]. However, accurately annotating multi-class cell arrangements in pathology images is challenging, as distinct cell types display unique spatial patterns, requiring significant domain expertise. Although annotated datasets for multiple cell types exist, they often lack the diversity needed for generalization across various tissues and organs.
To alleviate the burden of manual annotation and enhance the efficiency of analysis, there has been growing interest in utilizing generative models. Early works use Generative Adversarial Networks (GANs) [20] for automatic generation of pathology images [2, 34, 10]. In recent years, diffusion models [61, 31, 50, 11, 73, 74, 45] have emerged as much more reliable alternatives, generating accurate, high-resolution histopathology images [71, 51, 48, 23, 4]. However, all these diffusion models are only trained to directly generate the histopathology images. Despite the impressive visual results, these powerful models deliver limited insight into the underlying biology. It is very hard to connect the learned distributions with human knowledge about tumor microenvionment. This makes it challenging to validate, generalize, or control these models.
We argue that a key issue is the lack of explicit generation of cells and their spatial arrangement. The spatial organization of cells and the interactions across different cell types are critical for understanding tumor microenvironments, disease progression, and tissue regeneration [59]. The density and spatial distribution of various cell types -such as lymphocytes, epithelial cells, and stromal cells- are essential for pathologists in making accurate diagnoses and prognoses. For instance, the detection and quantification of tumor-infiltrating lymphocytes (TILs), which are lymphocytes located within the boundaries of invasive tumors [56], have been strongly linked to improved clinical outcomes [63, 57]. The presence of isolated or small clusters of tumor cells at the invasive front, known as tumor budding, serves as a prognostic biomarker associated with a higher risk of lymph node metastasis in colorectal carcinoma and other solid malignancies [47].
In this paper, we explore the problem of generating cell spatial layouts, with multiple benefits. First and foremost, directly modeling cell layouts brings us closer to aligning with pathologists’ domain knowledge. This enables the direct verification of synthetic data through quantification and comparison with expert knowledge, facilitating greater accuracy and trust in the generated data. Additionally, it opens up the possibility of controlling the layout generation process, allowing the model to generalize to previously unseen scenarios. From a data augmentation perspective, generating cell layouts also enables the creation of histology images conditioned on these layouts. This capability allows for the production of synthetic images with cell annotations, which can significantly aid in the training of models for various downstream tasks, particularly cell detection and classification.
Figure 1: Illustrations of intra-class distribution and the inter-class relationship across various cell types. Here we only highlight the tumor/epithelial and lymphocytes.
Cell topology is the key. To generate accurate cell layouts, we argue that the key is to model the topological relationships between cells, defined by their spatial organization, clustering, mixing, and connectivity. These topological patterns provide valuable insights into cellular communication, structural changes, and morphological abnormalities linked to pathological conditions such as cancer, inflammation, and fibrosis. Fig. 1 illustrates the significance of these topological relationships in understanding tissue organization and disease progression. These relationships reveal how cells are spatially arranged to form structural motifs that are crucial for tissue function and stability. For example, on the left, we show that in healthy glandular tissues, epithelial cells typically organize into circular or tubular structures, forming acinar units that are essential for secretion and nutrient transport [59]. In tissues with diverse cell types, topological relationships capture spatial configurations and connectivity patterns that reveal cellular interactions critical for tissue health or disease [72]. For example, immune cells clustering around tumor cells (Fig. 1 (right)) or specific fibroblast-epithelial arrangements highlight processes like immune surveillance, inflammation, and stromal support. By identifying these patterns, topological methods offer insights into multi-class cellular interactions, revealing potential biomarkers and enhancing our understanding of disease progression.
We propose the first diffusion model that generates cell topology for digital pathology. Our method guides the generation process with both 0- and 1-dimensional topological features, i.e., cell clusters and holes/gaps enclosed by cells. We implement this using the theory of persistent homology, which models cell topology in a multi-scale manner. As illustrated in Fig. 1, learning the topology ensures that the generated cell layouts not only preserve intra-class spatial/structural property but also accurately capture inter-class interactions.
The second contribution of this paper is a novel cell counting loss that ensures the diffusion model learns the correct distribution of cell numbers from the data. This addresses a key challenge in previous diffusion models for histopathology images [23, 4, 44], which is often biased toward unrealistically low or high cell counts.
By ensuring both accurate cell topology and cell count, our method generates layouts that more closely resemble real tissue microenvironments. This precision in modeling cell distributions and interactions enhances the utility of generated images for augmenting training in downstream tasks such as cell detection and classification, as demonstrated in our experiments. From a modeling perspective, our approach moves closer to generating biologically faithful representations of cellular environments, integrating human domain knowledge, and improving generalization to diverse scenarios.
Finally, we introduce the Topological Fréchet Distance (TopoFD), a novel metric designed to assess the spatial and topological accuracy of the generated layouts. TopoFD measures the similarity between the topological features of real and synthetic cell layouts, whereas traditional metrics such as FID focus solely on visual similarity.
In summary, our contribution is three-fold:
  • We present the first diffusion model designed to generate cell topology for digital pathology, simulating realistic intra- and inter-class spatial distributions in the generated layouts.
  • We introduce a novel cell counting loss that aligns the generated cell numbers with real data, addressing biases in prior diffusion models and ensuring realistic cell density in synthetic layouts.
  • We introduce the Topological Fréchet Distance (TopoFD), a new metric designed to evaluate the topological similarity of generated cell layouts.
Extensive experiments demonstrate the effectiveness of our proposed method, showing that it not only enhances sample quality but also significantly improves performance in downstream tasks such as cell detection and classification.
2 Related Works
Diffusion Models for Digital Pathology. Diffusion models, such as Denoising Diffusion Probabilistic Models (DDPMs) [31, 61, 50, 62, 11, 75, 15, 33] and Latent Diffusion Models (LDMs) [53, 55, 6], have significantly advanced image synthesis by modeling data distributions through iterative denoising processes on data sample or latent spaces. These models have been adapted to generate high-fidelity histopathology images in digital pathology [71, 4, 23, 28]. However, these methods often overlook the significance of multi-class cell layouts, which are crucial for accurately representing tissue structures.
Recent approaches to synthetic cell layout generation in digital pathology have explored nuclei labeling and cell arrangement. Abousamra et al. utilized spatial statistics and topological descriptors in GANs to model complex cell configurations [2]. DiffMix employs conditioned diffusion models for augmenting imbalanced nuclear pathology datasets, generating more realistic images than earlier methods [51]. Similarly, a inter diffusion framework proposed in [48] generates paired histopathology images and nuclei labels simultaneously, enhancing the context-awareness of synthetic data. Another strategy, a diffusion model focused on cell layout generation, uses density maps to incorporate spatial distributions, though it simplifies cell counts into five categorical conditions, which provides only limited guidance [44]. While these models have contributed to more realistic cell layout synthesis, they generally lack fine control over cell density and fail to explicitly preserve intra-class spatial distributions and inter-class spatial relationships across cell types. This limits their effectiveness in replicating the complex and varied structures of real tissue.
Topology-Driven Methods in Deep Learning. Incorporating algebraic topology [49] in deep learning frameworks is becoming increasingly significant. A key development in this front has been persistent homology [13, 14], a mathematical theory that analyzes how structural features (such as connected components, loops, and voids) persist across different scales in data. This approach has proven particularly valuable for understanding complex data structures while being resistant to noise. Its applications range from image segmentation [36, 8, 70, 64], feature extraction [19] and disease diagnosis [67, 43, 69, 42]. The field has further expanded through additional mathematical frameworks, including discrete Morse theory [17, 37, 38], homotopy warping [35], structural relationship studies [25, 27], and methods for analyzing shape features [60, 68].
For generative models, topological information has become increasingly essential, particularly in biomedical imaging, where capturing both visual realism and topological integrity is critical. One notable method is TopoGAN [66], which integrates topology into GANs by introducing a novel topological loss based on persistent homology. Recently, TopoDiffusionNet [26] has integrated topology with diffusion models, introducing a topology-based objective function that guides the model’s denoising process to generate images with exact object counts. However, these methods, designed for natural images, are unsuitable for our cell topology generation task. Their emphasis on overall structure numbers cannot capture the cell structural pattern and cell spatial interactions among multiple cell types as our method does.
Figure 2: An overview of our method TopoCellGen. (a) denotes the overview workflow. (b) shows the details of $\mathcal{L}_{\text{count}}$, $\mathcal{L}_{\text{intra}}$ and $\mathcal{L}_{\text{inter}}$.
3 Method
In this section, we introduce our methodology for generating realistic multi-class cell layouts with precise topological control. To accurately control the cell density in the generated layout, we employ a conditioning mechanism based on the cell count for each type. In the multi-class setting, assuming we have $n$ classes of cells, we define the condition vector $c=[c_{1},c_{2},...,c_{n}]$, where each element $c_{i}$ represents the count of cells in the respective channel.
However, the condition vector $c$ alone is insufficient to precisely control cell counts [26]. Additionally, it cannot preserve the intra-class spatial distributions as well as inter-class topological relationships in the generated layouts. To address this, we formulate a cell counting loss to strengthen the control on accurate cell counts and introduce two topology-aware objective functions based on persistent homology [13, 49] to guide the generation process, ensuring realistic and contextually accurate cell layouts that respect both spatial and topological constraints.
The remainder of this section is structured as follows: In Sec. 3.1, we present a brief introduction to diffusion models. Next, in Sec. 3.2, we provide a quick background on persistent homology. Following this, we introduce TopoCellGen, our topology-preserving cell layout generation method, in Sec. 3.3. We introduce our proposed Topological Fréchet Distance (TopoFD) in Sec. 3.4. Finally, in Sec. 3.5, we present the overall inference pipeline and how the generated cell layouts are transformed into H&E images for training augmentation for downstream tasks.
Preliminaries. In a multi-class cell layout containing $n$ channels, each channel corresponds to a specific cell type (e.g., lymphocyte), with each cell represented as a square where the pixel value is set to 1, while the pixel value of the background is set to 00.
3.1 Diffusion Models
Our generative approach utilizes a denoising diffusion probabilistic model (DDPM) [31], which learns to reverse a forward process that incrementally adds Gaussian noise to transform a structured cell layout into a noise distribution. The reverse process then reconstructs the layout from noise via iterative denoising.
Let $x_{0}$ represent the target cell layout and $x_{T}$ denote pure Gaussian noise. At each time step $t$, noise is added to the data based on a variance schedule $\beta_{t}$: $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$. This forward process results in the progressively noisier version of the data, with $x_{T}$ approximating an isotropic Gaussian distribution as $t$ increases.
The reverse process, parameterized by a neural network $\epsilon_{\theta}(x_{t},t)$ (typically UNet [54]), learns to iteratively denoise $x_{T}$ back to $x_{0}$. Conditioning on the cell count vector $c$, the model is trained by minimizing a simplified variant of the variational lower bound, specifically focusing on predicting the noise added at each step:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},c,t)\|^{2}\right]$ (1)
where $\epsilon\sim\mathcal{N}(0,I)$ is the noise sampled during training. This objective enables the model to learn the reverse process effectively. Instead of the standard iterative denoising, we also approximate the noiseless layout $\widehat{x}_{0}^{t}$ deterministically for any noisy state $x_{t}$ by marginalizing over the noise schedule:
$\widehat{x}_{0}^{t}\approx\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},c,t)\right)\vspace{-.1in}$ (2)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, which aggregates the effect of the variance schedule up to time $t$. This predicted noiseless layout, $\widehat{x}_{0}^{t}$, will be used to impose constraints in subsequent stages.
3.2 Background: Persistent Homology
In algebraic topology [49], homology classes provide a structured way to capture topological features across multiple dimensions. For instance, 0-, 1-, and 2-dimensional features represent connected components, loops, and voids, respectively. The $d$-dimensional Betti number, $\beta_{d}$, quantifies the number of $d$-dimensional features present, offering insights into the underlying topological complexity. However, extending these concepts to real-world data, which is often continuous and noisy, introduces challenges for accurately capturing topological structures.
Persistent homology, developed in the early 2000s [13, 14], addresses this need by tracking the evolution of topological features across multiple scales, making it particularly effective for discrete datasets like cell point clouds. The process begins by constructing a filtration—a sequence of nested simplicial complexes built from the point cloud data by incrementally connecting points based on a scale parameter. As the parameter varies, topological features such as connected components and loops appear and eventually vanish, each represented by a point in a persistence diagram (Dgm). In a diagram, each point $(b,d)$ marks the birth and death of a feature, capturing its persistence across scales and providing a compact, multi-scale summary of the underlying topology. More details are in the Supplementary.
3.3 Spatially Aligned Cell Layout Generation
The primary objective of our method is to generate multi-class cell layouts that accurately simulate both the topological and spatial properties of real-world biological cell distributions. To achieve this, we ensure accurate cell counts for each cell type through a cell counting loss, while also preserving spatial relationships within individual cell types via enforcing intra-class spatial consistency. Furthermore, we maintain structural coherence across all cell types by applying an inter-class structural regularization, leveraging 1-dimensional persistent homology to encapsulate both type-specific and collective spatial properties. The overall pipeline is shown in Fig. 2.
Cell Counting Loss. Given the target layout $x_{0}$, which serves as the ground truth, for each time step $t$, we obtain the predicted noiseless layout $\widehat{x}_{0}^{t}$ using Eq. 2. To ensure precise control over the number of cells in the generated layout, we introduce a differentiable cell counting loss. The key challenge lies in making the counting operation differentiable for gradient-based optimization. We address this by employing the Straight-Through Estimator (STE) [5], which enables gradient flow through the discrete binarization operation. Specifically, after obtaining $\widehat{x}_{0}^{t}$, we apply a hard threshold to obtain binary values:
$b(\widehat{x}_{0}^{t})=float((\widehat{x}_{0}^{t}\geq\tau))$ (3)
where $\tau$ is the threshold parameter. Here we set it to the median value of $\widehat{x}_{0}^{t}$. During back-propagation, the STE treats the thresholding operation as an identity function, allowing gradients to flow through. The cell counting loss is then formulated as:
$\mathcal{L}_{\text{count}}=\frac{1}{|n|}\sum_{i=1}^{n}\left|\frac{\sum b(\widehat{x}_{0}^{t})^{(i)}}{\delta}-\frac{\sum x_{0}^{(i)}}{\delta}\right|\vspace{-.1in}$ (4)
where $b(\widehat{x}_{0}^{t})^{(i)}$ represents the binarized prediction for the $i$-th channel, and $\delta$ indicates the area ($3\times 3$) of a single cell in the layouts. This formulation provides a differentiable approximation to the discrete cell counting operation, enabling end-to-end training while maintaining precise control over the number of cells for each cell type.
Intra-Class Spatial Consistency. To enforce spatial consistency within each cell type, we first calculate the distance transform map [16] for each channel in both the target layout $x_{0}$ and the predicted noiseless layout $\widehat{x}_{0}^{t}$. The distance transform $D(x)$ is a function that assigns to each pixel the minimum Euclidean distance to the nearest cell (or non-zero pixel) in the channel. This can be formally written as:
$D(x)=\min_{p\in\text{cells}}\|x-p\|\vspace{-.1in}$ (5)
where $p$ represents the positions of cells in the layout. After obtaining the distance transform maps of the target layout and the predicted noiseless layout, $\widehat{x}^{edt}_{t}=D(b(\widehat{x}_{0}^{t}))$ and $x^{edt}_{0}=D(x_{0})$, we calculate the 1-dim persistence diagrams for both of them, $Dgm(\widehat{x}^{edt}_{t})$ and $Dgm(x^{edt}_{0})$ respectively. Similar to previous topological losses [36], we will use the classic Wasserstein distance between the two diagrams. Given two diagrams $Dgm(q)$ and $Dgm(s)$, the $p$-th Wasserstein distance is defined as follows:
$W_{p}(Dgm(q),Dgm(s))=\left(\underset{\gamma\in\Gamma}{inf}\sum\limits_{x\in Dgm(q)}{||x-\gamma(x)||}^{p}\right)^{\frac{1}{p}}\vspace{-.1in}$
where $\Gamma$ represents all bijections from $Dgm(q)$ to $Dgm(s)$.
The Wasserstein distance operates by identifying an optimal correspondence between points in two diagrams, assigning unmatched points to their projections on the diagonal. This distance metric is calculated by summing the distances between all paired points. The process of finding this optimal matching, as well as calculating the Wasserstein distance, can be accomplished using either the traditional Hungarian algorithm or more sophisticated methods [39, 41].
Next, we denote $\gamma^{\ast}$, the optimal matching between $Dgm(\widehat{x}^{edt}_{t})$ and $Dgm(x^{edt}_{0})$. Each persistence dot in $Dgm(\widehat{x}^{edt}_{t})$ is matched either to a target dot in $Dgm(x^{edt}_{0})$ or its projection on the diagonal. We can now formulate the spatial distribution consistency loss as the squared distance between every dot in $Dgm(\widehat{x}^{edt}_{t})$ and its match:
$\mathcal{L}_{\text{spc}}=\sum\limits_{q\in Dgm(\widehat{x}^{edt}_{t})}||q-\gamma^{\ast}(q)||^{2}\vspace{-.1in}$ (6)
For a multi-class cell layout containing $n$ classes of cells, we formulate the intra-class spatial consistency loss as follows by averaging the Eq. 6 across multiple classes:
$\mathcal{L}_{\text{intra}}=\frac{1}{|n|}\sum_{i=1}^{n}\mathcal{L}_{\text{spc}}\left(Dgm((\widehat{x}^{edt}_{t})^{(i)})),Dgm((x^{edt}_{0})^{(i)})\right)$ (7)
Inter-Class Structural Regularization. Beyond maintaining spatial distribution consistency within individual cell types, it is equally important to capture the relationships between different cell types. To achieve this, we construct a unified layout by combining all cell types into a single-channel representation, referred to as the aggregated layout: $x_{0}^{agg}=Agg(x_{0})$ and $\widehat{x}_{0}^{t,agg}=Agg(\widehat{x}_{0}^{t})$. We then compute the distance transform for the aggregated layouts, with $\widehat{x}^{edt}_{t,agg}=D(\widehat{x}_{0}^{t,agg})$ representing the distance transform of the predicted layout and $x^{edt}_{0,agg}=D(x_{0}^{agg})$ for the target layout. The inter-class structural loss $\mathcal{L}_{\text{inter}}$ is computed similarly to the intra-class loss:
$\mathcal{L}_{\text{inter}}=\mathcal{L}_{\text{spc}}\left(Dgm(\widehat{x}^{edt}_{t,agg}),Dgm(x^{edt}_{0,agg})\right)\vspace{-.1in}$ (8)
Together, these class-specific and cross-class regularizations ensure that both individual cell distributions and their cumulative spatial interactions are enforced, preserving critical spatial dynamics within and between cell types in the generated layouts.
Final Objectives. The final training objective function of the model is the weighted sum of the three losses with $\mathcal{L}_{\text{simple}}$:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{simple}}+\lambda_{c}\mathcal{L}_{\text{count}}+\lambda_{\text{intra}}\mathcal{L}_{\text{intra}}+\lambda_{\text{inter}}\mathcal{L}_{\text{inter}}$ (9)
where $\lambda_{c}$, $\lambda_{\text{intra}}$ and $\lambda_{\text{inter}}$ are hyper-parameters that control the relative contributions of the respective loss terms.
3.4 Topological Fréchet Distance (TopoFD)
Fréchet Inception Distance (FID) [30] is a key metric for evaluating the quality of generated image data by comparing the distributions of real and generated samples. FID operates by computing the mean and covariance of feature representations extracted from a pre-trained model, typically InceptionV3 [65]. The distance between these distributions is then measured using the Fréchet distance [18]. The formula for FID is as follows:
$\text{FID}=\|\mu_{r}-\mu_{g}\|^{2}+\text{Tr}(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{\frac{1}{2}})$ (10)
where $\mu_{r}$, $\Sigma_{r}$ and $\mu_{g}$, $\Sigma_{g}$ are the mean and covariance of the real and generated data, respectively.
Figure 3: Intuition of our proposed Topological Fréchet Distance. TCE indicates the Total Count Error.
Traditional FID measures feature similarities but fail to capture complex spatial and topological cell interactions. For example, in Fig. 3, synthetic layouts of three cell types (lymphocyte, tumor/epithelial and stromal cells) demonstrate this limitation. Both Set 1 and Set 2 yield similar FID scores with respect to the Ref Set, yet differ significantly from each other in spatial configuration. Set 1 mirrors the reference with cohesive single cycles per cell type, while Set 2 presents multiple distinct cycles, deviating from the reference. The Total Count Error (TCE) reveals identical counts between Set 1 and the reference (TCE = 0) but substantial discrepancies for Set 2 (TCE = 10). Despite matching counts, Set 1’s spatial arrangement still diverges from the reference. This suggests that a count metric alone does not ensure spatial and topological fidelity. We propose Topological Fréchet Distance (TopoFD) to capture higher-dimensional topological features, enabling greater sensitivity to spatial configurations essential in synthetic pathology data generation.
The pipeline of our Topological Fréchet Distance (TopoFD) is shown in Fig. 4, taking one type of cell as an example. For each layout in the reference and synthetic sets, we first obtain a point cloud, with the points corresponding to the cell center coordinates. Then, we calculate 1-dimensional persistence diagrams of each point cloud in each set:
$Dgm_{ref}=\{Dgm_{r}^{1},Dgm_{r}^{2},...,Dgm_{r}^{n}\}$
$Dgm_{syn}=\{Dgm_{s}^{1},Dgm_{s}^{2},...,Dgm_{s}^{n}\}$
where $Dgm_{r}^{i}$ and $Dgm_{s}^{j}$ are persistence diagrams in reference and synthetic sets, $n$ is the number of the samples. We then compute each set’s barycenter [9, 7], which minimizes the sum of Wasserstein distances [9] between individual diagrams and the barycenter. For the reference set, the barycenter $\overline{Dgm}_{real}$ is given by:
$\overline{Dgm}_{r}=\arg\min_{\overline{Dgm}_{r}}\sum_{i=1}^{n}W_{2}^{2}(Dgm_{r}^{i},\overline{Dgm}_{r})$ (11)
where $W_{2}$ is the 2-Wasserstein distance. A similar process is applied for $\overline{Dgm}_{s}$ from the synthetic set.
Figure 4: The overall pipeline of calculating the Topological Fréchet Distance. Take the lymphocyte as an example.
Figure 5: The qualitative results of our proposed TopoCellGen. Columns $1$-$3$: BRCA-M2C dataset. Columns $4$-$6$: Lizard dataset. The cell types and their corresponding colors are shown on the right side of the figure.
To enable statistical analysis, we transform the persistence diagrams and their barycenters into persistence landscapes, transforming each diagram into a sequence of continuous, piecewise linear functions that preserve topological information while enabling standard statistical operations [7]. The vectorized barycenters are regarded as mean vectors $\boldsymbol{\mu}$ and we compute the covariance matrices $\boldsymbol{\Sigma}$ from both sets of persistence diagrams.
For multi-class layouts, TopoFD is calculated in two steps. First, for each cell type, we compute the vectorized persistence diagrams for both reference and synthetic layouts containing that type. Then, the final TopoFD is the average Fréchet Distance across cell types:
$\text{TopoFD}=\frac{1}{|n|}\sum_{i=1}^{n}\text{FD}(\boldsymbol{\mu}_{i}^{r},\boldsymbol{\Sigma}_{i}^{r},\boldsymbol{\mu}_{i}^{s},\boldsymbol{\Sigma}_{i}^{s})$ (12)
where $\text{FD}(\cdot)$ denotes the Fréchet distance computation.
3.5 Layout-Guided Pathology Image Generation
In the final stage of our methodology, we focus on generating high-resolution pathology images based on the previously created cell layouts. These layouts serve as spatial and structural guides, ensuring that the pathology images adhere to the organization of cells within the layout. We employ a guided diffusion model [11] to synthesize realistic pathology images, conditioned on the underlying cell arrangement. This approach ensures that cells’ spatial distribution and organization in the generated images align closely with the provided cell layouts. More details are provided in the Supplementary.
4 Experiments
We conduct extensive experiments on two public and widely used nuclei analysis datasets. We compare our method against SoTA layout generation methods in terms of sample quality and performance on downstream tasks.
Datasets. We evaluate our proposed method on TCGA Breast Cancer Cell Classification Dataset (BRCA-M2C) [1] and Lizard dataset [22]. More details about the datasets, the test set split, and the cell count distribution analysis are provided in the Supplementary.
Evaluation Metrics. We evaluate our proposed method on both sample quality and the performance of downstream tasks. We use Fréchet Inception Distance (FID) [30], the cell count error for each cell type, the total count error (TCE), our proposed TopoFD and maximum mean discrepancy (MMD) [66] to evaluate how well the generated cell layouts align with the reference layouts. Note that for FID, feature extraction is tailored to each dataset with custom-trained models. On the other hand, we generate $2,000$ image-layout pairs as augmented training data for cell detection and classification tasks, evaluating their performance with the F1-score. More details are provided in the Supplementary.
Implementation Details are in the Supplementary.
Backpropagation from Distance Transform Map to Binary Mask. To backpropagate the gradient from the distance transform map to a binary mask, we leverage an efficient implementation of the geodesic distance transform provided by the FastGeodis library [3]. This method enables the computation of a differentiable distance map, which allows us to propagate gradients through the distance transform during training. The library implements a parallelized raster scan method that efficiently computes the Euclidean distance transforms on GPU hardware.
4.1 Experimental Results
Qualitative Results. The qualitative results in Fig. 5 illustrate that our proposed method, TopoCellGen, successfully generates cell layouts that closely align with the reference layouts in both cell density and spatial arrangement. In each example, the generated layouts demonstrate a high degree of consistency in cell distribution, preserving the density patterns observed in the reference cell layouts. Furthermore, the spatial relationships between cell types are maintained, with distinct inter-class interactions that reflect the structural characteristics of the reference layouts.
Quantitative Results on Sample Quality.  Tab. 1 shows the results of sample quality and generation accuracy on the BRCA-M2C and Lizard datasets compared with the state-of-the-art cell layout generation baselines: ADM  [11], TMCCG  [2] and Spatial Diffusion  [44]. Across all multi-class datasets, our proposed TopoCellGen demonstrates superior performance by achieving the lowest FID and TopoFD. This means that the generated samples of TopoCellGen have the highest visual fidelity and topological accuracy, closely approximating real multi-class cell distributions. Beyond this, our TopoCellGen also greatly reduces the count errors across multiple cell classes, resulting in more accurate and topologically consistent synthetic cell layouts compared to other methods. These results emphasize the capability of TopoCellGen to maintain realistic cell distribution and topological relationships, meanwhile achieving precise control of cell density.
Method FID $\downarrow$ Lym. $\downarrow$ Epi. $\downarrow$ Stro. $\downarrow$ Neu. $\downarrow$ Pla. $\downarrow$ Eos. $\downarrow$ Con. $\downarrow$ TCE $\downarrow$ TopoFD $\downarrow$ MMD $\downarrow$
BRCA-M2C ADM [11] 1.150 13.757 40.230 15.491 22.465 133.012 0.732
TMCCG [2] 0.634 11.503 34.032 12.907 19.687 89.252 0.635
Spatial Diffusion [44] 0.263 10.852 35.954 13.496 20.806 97.584 0.589
TopoCellGen 0.005 2.090 3.824 2.468 5.192 69.354 0.421
Lizard ADM [11] 0.059 16.508 11.796 1.123 4.328 1.598 10.737 23.964 65.910 0.783
TMCCG [2] 1.093 15.548 10.011 2.376 4.293 1.872 11.643 22.604 63.120 0.667
Spatial Diffusion [44] 0.137 10.740 9.062 3.040 6.552 2.173 11.225 20.606 79.591 0.883
TopoCellGen 0.027 6.155 6.560 1.022 2.982 1.167 7.288 11.590 31.607 0.536
Table 1: Results for BRCA-M2C and Lizard datasets on the quality of the generated samples.
Data Method F1-Score $\uparrow$
Lymphocytes Epithelial Stromal Mean Detection
Real. UNet $0.569\pm 0.010$ $0.736\pm 0.012$ $0.507\pm 0.015$ $0.604\pm 0.011$ $0.857\pm 0.006$
Real+Syn. (Rand) $0.549\pm 0.009$ $0.693\pm 0.014$ $0.472\pm 0.016$ $0.571\pm 0.013$ $0.848\pm 0.008$
Real+Syn (TMCCG) $0.650\pm 0.007$ $0.768\pm 0.010$ $0.511\pm 0.012$ $0.643\pm 0.009$ $0.852\pm 0.005$
Real+Syn (SpaDM) $0.647\pm 0.006$ $0.797\pm 0.003$ $0.554\pm 0.011$ $0.666\pm 0.007$ $0.853\pm 0.005$
Real+Syn (TopoCellGen) $\boldsymbol{0.656\pm 0.003}$ $\boldsymbol{0.803\pm 0.005}$ $\boldsymbol{0.574\pm 0.004}$ $\boldsymbol{0.678\pm 0.004}$ $\boldsymbol{0.860\pm 0.004}$
Real. MCSpatNet $0.615\pm 0.008$ $0.777\pm 0.010$ $0.540\pm 0.013$ $0.644\pm 0.009$ $0.855\pm 0.005$
Real+Syn. (Rand) $0.578\pm 0.009$ $0.756\pm 0.012$ $0.502\pm 0.014$ $0.612\pm 0.010$ $0.851\pm 0.006$
Real+Syn (TMCCG) $\boldsymbol{0.678\pm 0.006}$ $0.800\pm 0.005$ $0.522\pm 0.014$ $0.667\pm 0.007$ $0.853\pm 0.004$
Real+Syn (SpaDM) $0.639\pm 0.005$ $0.804\pm 0.007$ $0.563\pm 0.012$ $0.669\pm 0.006$ $0.855\pm 0.005$
Real+Syn (TopoCellGen) $0.652\pm 0.004$ $\boldsymbol{0.817\pm 0.006}$ $\boldsymbol{0.582\pm 0.005}$ $\boldsymbol{0.684\pm 0.004}$ $\boldsymbol{0.862\pm 0.004}$
Table 2: Results on cell detection and classification tasks on BRCA-M2C dataset. The best and statistically significant results are highlighted in bold.
Performance on Downstream Tasks. In Tab. 2, we show the results of cell detection and classification tasks using our synthetic image-layout pairs as data augmentations. We present the results using two frameworks, UNet [54] and MCSpatNet [1]. The results indicate that TopoCellGen achieves the highest F1 scores across various cell types, including inflammation and epithelial cells, resulting in superior mean F1 scores and detection metrics on the BRCA-M2C dataset. This demonstrates our TopoCellGen’s capability to capture both spatial fidelity and topological accuracy, ensuring that synthetic data closely resembles real biological structures. By accurately modeling complex spatial distributions and inter-class relationships, TopoCellGen provides biologically plausible synthetic samples that improve the generalizability of detection and classification models. Furthermore, its balanced representation of cellular compositions reduces class-wise biases, allowing the classifiers to better learn fine-grained distinctions.
4.2 Ablation Studies
Extensive experiments are conducted to elucidate the effectiveness and robustness of our loss components and hyper-parameters. All experiments are performed on the BRCA-M2C dataset.
Ablation Study on Loss Components. We evaluate the contributions of three loss functions to the model’s performance in generating synthetic cell layouts, measured by FID, Total Counting Error, and TopoFD. As shown in Tab. 3, the results demonstrate that $\mathcal{L}_{\text{count}}$ significantly reduces the counting errors, achieving better accuracy when combined with the other losses. The $\mathcal{L}_{\text{intra}}$ improves the fidelity and topological accuracy of the generated multi-class layouts, as indicated by lower FID and TopoFD. The $\mathcal{L}_{\text{inter}}$ has a smaller effect on the cell counting error but enhances structural consistency, as reflected in improved TopoFD. Using all three losses together yields the best results across all metrics, demonstrating their complementary roles in achieving accurate and realistic multi-class cell layouts.
$\mathcal{L}_{\text{count}}$ $\mathcal{L}_{\text{intra}}$ $\mathcal{L}_{\text{inter}}$ BRCA-M2C
FID $\downarrow$ TCE $\downarrow$ TopoFD $\downarrow$
1.150 22.465 133.012
0.842 12.253 118.304
0.621 16.315 98.798
1.083 21.928 126.742
0.232 6.854 85.672
0.498 8.012 91.324
0.327 17.573 73.612
0.005 5.192 69.354
Table 3: Ablation study on loss components.
Ablation Study on Loss Weights. The ablation study on loss weights is presented in Tab. 4. The results indicate that when the loss weights are higher, such as $0.005$ or $0.001$, they impose overly strong regularization, leading to suboptimal outcomes. Conversely, when the loss weights are lowered to $41e-$ 4 or even $55e-$ 5, the regularization becomes too weak to achieve optimal performance. The configuration with moderate weights $45e-$ 4 achieves the best balance, yielding the lowest FID, total count error, and TopoFD.
$\lambda_{c}$ $\lambda_{\text{intra}}$ $\lambda_{\text{inter}}$ BRCA-M2C
FID $\downarrow$ TCE $\downarrow$ TopoFD $\downarrow$
0.005 0.005 0.005 0.289 14.920 92.718
0.001 0.001 0.001 0.153 12.471 85.642
1e-4 1e-4 1e-4 0.012 8.275 88.627
5e-5 5e-5 5e-5 0.129 11.378 83.629
5e-4 5e-4 5e-4 0.005 5.192 69.354
Table 4: Ablation study on loss weights.
5 Conclusion
In conclusion, TopoCellGen presents a robust framework for generating realistic cell topologies in digital pathology by integrating 0- and 1-dimensional topological features within a diffusion model. This method addresses significant limitations of existing approaches, including lack of control over cell count and challenges in accurately capturing both intra-class spatial patterns and inter-class interactions. By integrating topological features into the generative process, our model preserves essential structural characteristics, resulting in realistic and contextually accurate cell layouts. Additionally, the proposed TopoFD metric provides a thorough evaluation of topological fidelity, validating our model’s effectiveness in representing complex spatial and topological relationships. Experimental results demonstrate that our TopoCellGen produces synthetic cell layouts that closely mimic real tissue structures and enhances performance in tasks like cell detection and classification, highlighting its potential to enrich training data and support advancements in digital pathology research.

### v2

@@
 Experimental results demonstrate the effectiveness of our approach in generating multi-class cell layouts that capture intricate topological relationships.
+Code is available at https://github.com/Melon-Xu/TopoCellGen.
 1 Introduction
-Deep-learning methods have made substantial advances in foundational tasks for nuclei analysis, including instance segmentation [29, 21, 40, 32], classification, and detection [1, 52].
+Deep-learning methods have made substantial advances in foundational tasks for nuclei analysis, including instance segmentation [29, 21, 40, 32, 73], classification, and detection [1, 52].
 These tasks provide the basis for downstream analyses, enabling detailed characterization of tissue architecture and cellular interactions, which are crucial for diagnostic and prognostic applications in pathology [46, 12].
@@
 Early works use Generative Adversarial Networks (GANs) [20] for automatic generation of pathology images [2, 34, 10].
-In recent years, diffusion models [61, 31, 50, 11, 73, 74, 45] have emerged as much more reliable alternatives, generating accurate, high-resolution histopathology images [71, 51, 48, 23, 4].
+In recent years, diffusion models [61, 31, 50, 11, 75, 76, 45, 69] have emerged as much more reliable alternatives, generating accurate, high-resolution histopathology images [72, 51, 48, 23, 4].
 However, all these diffusion models are only trained to directly generate the histopathology images.
@@
 This capability allows for the production of synthetic images with cell annotations, which can significantly aid in the training of models for various downstream tasks, particularly cell detection and classification.
-Figure 1: Illustrations of intra-class distribution and the inter-class relationship across various cell types. Here we only highlight the tumor/epithelial and lymphocytes.
+Figure 1: Illustrations of intra-class distribution and the inter-class relationship across various cell types. Here we only highlight the tumor/epithelial, lymphocytes, and stromal cells.
 Cell topology is the key.
@@
 Implementation Details are in the Supplementary.
+Biological Plausibility.
+To further validate the quality of our generated cell layouts, we conducted an evaluation involving a board-certified pathologist with 7+ years of experience.
+From a pathologist’s perspective, the synthetic cell layouts are nearly indistinguishable from real ones and accurately capture the defining benign or malignant features of their real counterparts.
+More details are in the Supplementary.
 Backpropagation from Distance Transform Map to Binary Mask.
@@
 Qualitative Results.
-The qualitative results in Fig.
-5 illustrate that our proposed method, TopoCellGen, successfully generates cell layouts that closely align with the reference layouts in both cell density and spatial arrangement.
-In each example, the generated layouts demonstrate a high degree of consistency in cell distribution, preserving the density patterns observed in the reference cell layouts.
-Furthermore, the spatial relationships between cell types are maintained, with distinct inter-class interactions that reflect the structural characteristics of the reference layouts.
+Qualitative results are shown in Fig.
+5.
+Our proposed TopoCellGen method demonstrates cell layouts that closely match the reference layouts in both density and spatial arrangement.
+The generated layouts exhibit a high degree of distribution consistency, preserving density patterns and spatial relationships among cell types while reflecting the structural characteristics of the reference layouts.
 Quantitative Results on Sample Quality.
@@
 1 shows the results of sample quality and generation accuracy on the BRCA-M2C and Lizard datasets compared with the state-of-the-art cell layout generation baselines: ADM  [11], TMCCG  [2] and Spatial Diffusion  [44].
-Across all multi-class datasets, our proposed TopoCellGen demonstrates superior performance by achieving the lowest FID and TopoFD.
-This means that the generated samples of TopoCellGen have the highest visual fidelity and topological accuracy, closely approximating real multi-class cell distributions.
-Beyond this, our TopoCellGen also greatly reduces the count errors across multiple cell classes, resulting in more accurate and topologically consistent synthetic cell layouts compared to other methods.
-These results emphasize the capability of TopoCellGen to maintain realistic cell distribution and topological relationships, meanwhile achieving precise control of cell density.
+Across all multi-class datasets, our proposed TopoCellGen attains the lowest FID and TopoFD, indicating superior visual fidelity and topological accuracy.
+Moreover, TopoCellGen notably reduces multi-class count errors, resulting in more precise and topologically consistent synthetic cell layouts.
+These findings highlight TopoCellGen’s capacity to preserve realistic cell distributions, maintain topological relationships, and enable accurate control over cell density.
 Method FID $\downarrow$ Lym.
@@
 The $\mathcal{L}_{\text{inter}}$ has a smaller effect on the cell counting error but enhances structural consistency, as reflected in improved TopoFD.
-Using all three losses together yields the best results across all metrics, demonstrating their complementary roles in achieving accurate and realistic multi-class cell layouts.
+Combining all three losses yields superior results across all metrics, highlighting their complementary contributions to accurate and realistic multi-class cell layouts.
 $\mathcal{L}_{\text{count}}$ $\mathcal{L}_{\text{intra}}$ $\mathcal{L}_{\text{inter}}$ BRCA-M2C
@@
 5 Conclusion
-In conclusion, TopoCellGen presents a robust framework for generating realistic cell topologies in digital pathology by integrating 0- and 1-dimensional topological features within a diffusion model.
-This method addresses significant limitations of existing approaches, including lack of control over cell count and challenges in accurately capturing both intra-class spatial patterns and inter-class interactions.
-By integrating topological features into the generative process, our model preserves essential structural characteristics, resulting in realistic and contextually accurate cell layouts.
-Additionally, the proposed TopoFD metric provides a thorough evaluation of topological fidelity, validating our model’s effectiveness in representing complex spatial and topological relationships.
-Experimental results demonstrate that our TopoCellGen produces synthetic cell layouts that closely mimic real tissue structures and enhances performance in tasks like cell detection and classification, highlighting its potential to enrich training data and support advancements in digital pathology research.
+In summary, TopoCellGen presents a robust framework for generating realistic cell topologies in digital pathology.
+It accurately preserves both intra- and inter-class spatial patterns, ensures cell count control, and achieves high structural fidelity.
+Experimental results confirm its close approximation of real tissue layouts, thereby enhancing downstream tasks such as cell detection and classification.
+Acknowledgment.
+We thank Dr.
+Michael L.
+Miller for evaluating the biological plausibility of our results.
+This research was partially supported by the National Science Foundation (NSF) grants CCF-2144901, IIS-2123920, IIS-2212046, the National Institute of Health (NIH) grants R01GM148970, R01GM148970-03S1, R21CA258493, R21CA258493-01A1, R01CA297843, and the Stony Brook Trustees Faculty Award.